# Can Many-Shot In-Context Learning Help Long-Context LLM Judges? See More, Judge Better!

### v1

@@ Abstract
Leveraging Large Language Models (LLMs) as judges for evaluating the performance of LLMs has recently garnered attention. Nonetheless, this type of approach concurrently introduces potential biases from LLMs, raising concerns about the reliability of the evaluation results. To mitigate this issue, we propose and study two versions of many-shot in-context prompts, Reinforced and Unsupervised ICL, for helping GPT-4o-as-a-Judge in single answer grading. Based on the designed prompts, we investigate the impact of scaling the number of in-context examples on the agreement and quality of the evaluation. Furthermore, we first reveal the symbol bias in GPT-4o-as-a-Judge for pairwise comparison and then propose a simple yet effective approach to mitigate it. Experimental results show that advanced long-context LLMs, such as GPT-4o, perform better in the many-shot regime than in the zero-shot regime. Meanwhile, the experimental results further verify the effectiveness of the symbol bias mitigation approach. The code and generated data are released in https://github.com/nick7nlp/SeeMoreJudgeBetter.
@@
Mingyang Song, Mao Zheng, Xuan Luo Tencent MLPD nickmysong@tencent.com
@@ 1 Introduction
Figure 1: Agreement between different versions of evaluation results by adopting GPT-4o as a zero-shot judge. $\hat{v}_{1}$ and $\hat{v}_{2}$ are the results based on Prompt(A) in Table 2. $v_{1}$, $v_{2}$, and $v_{3}$ are results based on Prompt(B) in Table 2. Prompts (A) and (B) differ in whether to output the rating first or later. The agreement evaluations show that Prompt (A) and (B) almost obtain the same agreement evaluation results, but the latter is convenient for constructing many-shot in-context examples, so we adopt the latter generated rationales in this study. $v_{1}$ vs. $v_{2}$ denotes comparing the first and second versions of evaluations. $v_{1}$ vs. $v_{2}$ vs. $v_{3}$ denotes the comparison between the three versions of evaluations.
Typically, when humans and LLMs grade answers involving model-generated chain-of-thought rationales, it is challenging to assign an exact score from 1-10, even if a reference correct answer is provided and each score has a particular description. Meanwhile, it is impossible to cover all situations in the descriptions of scores. Existing LLM-as-a-Judge approaches mainly adopt prompts with several aspects of constraints, which is undoubtedly very difficult because LLMs do not know exactly what kind of answer corresponds to each score.
Newly expanded context windows of LLMs allow researchers to investigate ICL with more shots than the few-shot regime, namely many-shot ICL. Agarwal et al. (2024) explore scaling in-context learning to hundreds or thousands of examples and find performance improvements across multiple tasks. More importantly, they show that using the many-shot in-context examples generated through the zero-shot ICL is effective, and the many-shot ICL may overcome pre-training biases of LLMs, whereas few-shot ICL struggles. Therefore, an intuitive idea is to use the many-shot ICL, allowing LLM-as-a-Judge to see the zero-shot evaluations of similar questions and answers first and then scoring examples before scoring. Therefore, an interesting question arises:
@@
Motivated by prior findings and the above issue, we verify the evaluation agreement of the widely used prompts of LLM-as-a-Judge Zheng et al. (2023), as shown in Table 2. Concretely, the agreement experiments are based on the entire test set of GSM8K Cobbe et al. (2021), where the inference answers are obtained based on LLaMA3-70B (The pipeline as illustrated as in Figure 2 and details in §$\S$ 3). Figure 1 presents that the agreement between the two versions of evaluations is low, with nearly half of the ratings being disagreement.
Inspired by the above phenomena and issues, in this paper, we investigate whether many-shot in-context learning helps long-context LLM judges. Motivated by Agarwal et al. (2024), we introduce two versions of many-shot ICL prompts via Reinforced ICL and Unsupervised ICL for LLM-as-a-Judge. The former utilizes in-context examples with model-generated rationales, and the latter removes rationales used in the former, details in §$\S$ A.1. Meanwhile, we reveal a novel symbol bias in GPT-4o-as-a-Judge and explore a simple approach for mitigating this issue. Experiments show that many-shot ICL can help GPT-4o-as-a-Jduge obtain higher quality and more agreement evaluation results. As the number of in-context examples increases, the quality and agreement of evaluation improves significantly. We further verify the effectiveness of the symbol bias mitigation approach.
@@
  • Firstly, we propose investigating many-shot ICL to help long-context LLM judges, such as GPT-4o. To the best of our knowledge, we are the first to introduce many-shot in-context learning to assist long-context LLMs as judges and scale the number of in-context examples to verify the effectiveness.
  • Second, inspired by the Counting-Stars benchmark Song et al. (2024), we reveal a novel potential bias in GPT-4o-as-a-Judge, namely symbol bias, and propose a simple yet effective approach to mitigate it.
  • Finally, through experimental results, many-shot ICL can help GPT-4o-as-a-Jduge obtain higher quality and more agreement evaluation results. Meanwhile, we demonstrate the effectiveness of the symbol bias mitigation method in this study.
@@
Figure 3: Agreement between two versions of evaluation results. Concretely, the bar corresponding to "0" on the x-axis represents the number of samples with equal and unequal ratings in comparing evaluation results obtained twice using GPT-4o as the judge in the zero-shot regime. In addition, the zero-shot generated rationales are used for Reinforced ICL and appended to the prompt for Unsupervised ICL. The bar corresponding to "$2^{n}$" on the x-axis represents the agreement evaluation of using the GPT-4o as a judge in the many-shot Reinforced ICL regime.
@@ 2.1 Background of Many-Shot ICL
LLMs excel at few-shot in-context learning, which involves learning from a few input-output examples (“shots”) provided in context at inference without weight updates Brown et al. (2020). Newly expanded long-context LLMs allow us to investigate ICL with hundreds of examples, namely the many-shot regime Agarwal et al. (2024). Due to some limitations, we only examine GPT-4o-as-a-Judge in a many-shot regime in this paper.
@@ 2.2 Recalling Judgment Biases
LLM judges usually possess certain biases, which have been widely explored Wang et al. (2023b); Wu and Aji (2023); Zheng et al. (2023); Chen et al. (2024). For example, positional bias refers to the phenomenon where, during pairwise comparisons, LLM judges tend to favor one side of a pair regardless of the actual quality of the answers. In this paper, we investigate leveraging many-shot ICL to help GPT-4o as a better Judge, maybe through reducing judgment biases.
@@ 2.3.1 Single Answer Grading
In this variation, the GPT-4o Judge is required to assign a score to a single answer directly. Inspired by the recent work Agarwal et al. (2024), we adopt the Reinforced ICL, which uses model-generated rationales as the in-context examples. Meanwhile, we also design the prompt within the Unsupervised ICL. The 4-shot Reinforced ICL and 1-shot Unsupervised ICL append 4-shots with prompts are presented in Table 1 and Table 5 (Appendix).
@@ 2.3.2 Pairwise Comparison
The GPT-4o Judge is presented with a question and two answers and tasked with determining which is better or declaring a tie. The prompt used for this scenario is given in Table 2 (Appendix).
Figure 4: Agreement between two versions of evaluation results.
Figure 5: Agreement between two versions of evaluation results.
@@ 3.1 Experimental Settings
To analyze the GPT-4o judge in many-shot regime, we adopt LLaMA3-70B to generate answers for each question in GSM8K Cobbe et al. (2021) with a temperature of 0.7. Details of the used dataset are provided in §$\S$ A.3 (Appendix). We use the training set of GSM8K as the sampling pool of in-context samples and use the first 200 samples of the test set in GSM8K for many-shot ICL experiments in this paper. Specifically, we randomly sample in-context examples for each ${K}$-shot prompt in each test data for reliable results. To ensure that using more shots provides additional information, any ${K}$-shot ICL prompt in our experiments includes all in-context examples from prompts with less than ${K}$ examples (all samples in the training set).
Furthermore, inspired by Counting-Stars Song et al. (2024), even if too many shots are provided, the GPT-4o may not be able to utilize all of them. Because, in the Counting-Stars benchmark, when the number of pieces of evidence reaches 32, long-context LLMs (e.g., GPT-4 Turbo, Gemini 1.5 Pro, and Claude3 Opus) may no longer accurately obtain all of them. Therefore, adding more many-shot in-context examples is probably not captured by LLMs for learning as a reference to judge. However, the many-shot regime and Counting-Stars are substantially different, so there is no noise from "haystack". Hence, we set the maximum number of the in-context examples to 128, i.e., 128-shot.
@@
3.2 Agreement Evaluation
We investigate the agreement between different versions of evaluation results generated by the GPT-4o judge. Here, the single answer rating can be converted into pairwise comparison results to evaluate the agreement. As shown in Figure 3, the bar corresponding to "$2^{n}$" on the x-axis represents the number of samples with equal and unequal ratings in comparing evaluation results obtained twice using GPT-4o as the judge in the many-shot Reinforced ICL regime. From the results, agreement improves as we increase the number of shots provided as in-context examples during inference.
By analyzing the previous studies Wang et al. (2023b); Zheng et al. (2023); Chen et al. (2024), we suppose it unnecessary to obtain utterly accurate evaluation results (because this is very difficult) by using LLMs as judges. It is only required to ensure that the evaluation results are highly agreed upon multiple times so that the pairwise comparison may be fair. From the results, we find that the many-shot ICL examples keep the evaluation of LLMs more in agreement, which is essential. We consider that the main reason may be that the many-shot in-context examples mitigate the potential bias of the GPT-4o judge in the zero-shot regime.
Meanwhile, we also implemented experiments in the many-shot Unsupervised ICL, but the results show that this regime may be unsuitable in the scenario of acting as a judge because the problem of each sample is the same, but the scored questions and answers are different. From the results in Figure 4 and Figure 5, it can be seen that agreement is related to the number of appended ICL samples but not to the number of unsupervised samples.
@@ 3.3 Quality Evaluation
After the agreement evaluation, an important question arises about: Does a high agreement refer to high-quality evaluations? Around this issue, we pair-wise compare evaluation ratings between zero-shot and many-shot regimes adopting the designed prompts shown in Table 3.
From Figure 7 Compare(A, B), we can find that the evaluations obtained in the many-shot regime are significantly better than those in the zero-shot regime. However, as mentioned before, the positional bias of GPT-4o-as-a-Judge may cause these results, so we performed a second comparison by swapping the positions. As shown in Figure 7 Compare(B, A), we observe that as in-context examples increase, the results in the many-shot regime gradually turn around, that is, a higher winning rate.
Revealing Symbol Bias. Inspired by Song et al. (2024), some LLMs revealed a preference for specific symbols under stress testing (details are described in the Appendix). Therefore, we propose a conjecture: Does the GPT-4o prefer to choose A rather than B? To verify this conjecture, we swap the answers corresponding to symbols A and B, as shown in Figure 3. From Figure 6 Compare(A, B) and Compare(B, A), it can be seen that the results are different from the above experiments. Actually, the results of Compare(A, B) and Compare(A, B) should be similar, and the results of Compare(B, A) and Compare(B, A) should be similar. This phenomenon shows that symbol bias does exist when adopting GPT-4o-as-a-Judge.
@@ 4 Related Work
LLMs have exhibited remarkable general generation capabilities, positioning themselves as powerful assistants Zhao et al. (2023); OpenAI (2023). With the rapid progression of LLMs, evaluating their proficiency in adhering to human instructions is imperative. Given the advanced capabilities of LLMs, researchers have begun adopting these models to judge the performance of LLMs in following human instructions Koo et al. (2023); Liusie et al. (2023); Liu et al. (2023); Zhu et al. (2023); Lu et al. (2023); Fu et al. (2023b); Zheng et al. (2023); Wang et al. (2023b); Chen et al. (2024). Notably, the evaluation paradigm introduced by Zheng et al. (2023) has gained widespread adoption. However, LLMs as judges are revealed to have potential biases Wang et al. (2023b); Chen et al. (2024), which leads to higher uncertainty and disagreement during the evaluation using LLMs, questioning the validity of LLM-as-a-Judge.

### v3

@@ Abstract
Leveraging Large Language Models (LLMs) as judges for judging the performance of LLMs has recently garnered attention. However, this type of approach is affected by the potential biases in LLMs, raising concerns about the reliability of the evaluation results. To mitigate this issue, we propose and study two versions of many-shot in-context prompts, which rely on two existing settings of many-shot ICL for helping GPT-4o-as-a-Judge in single answer grading to mitigate the potential biases in LLMs, Reinforced ICL and Unsupervised ICL. Concretely, the former utilizes in-context examples with model-generated rationales, and the latter without. Based on the designed prompts, we investigate the impact of scaling the number of in-context examples on the consistency and quality of the judgment results. Furthermore, we reveal the symbol bias hidden in the pairwise comparison of GPT-4o-as-a-Judge and propose a simple yet effective approach to mitigate it. Experimental results show that advanced long-context LLMs, such as GPT-4o, perform better in the many-shot regime than in the zero-shot regime. Meanwhile, the experimental results further verify the effectiveness of the symbol bias mitigation approach. The code and data are released in https://github.com/nick7nlp/SeeMoreJudgeBetter.
@@
Mingyang Song, Mao Zheng, Xuan Luo Tencent Hunyuan nickmysong@tencent.com
@@ 1 Introduction
Figure 1: Consistency between different versions of judgment results by adopting GPT-4o as a zero-shot judge. $\hat{v}_{1}$ and $\hat{v}_{2}$ are the results based on Prompt(A) in Table 3. $v_{1}$, $v_{2}$, and $v_{3}$ are results based on Prompt(B) in Table 3. Prompts (A) and (B) differ in whether to output the rating first or later. The consistency evaluations show that Prompt (A) and (B) almost obtain the agreement results, but the latter is convenient for constructing many-shot in-context examples, so we adopt the latter generated rationales in this study. $v_{1}$ vs. $v_{2}$ denotes comparing the first and second versions of evaluations. $v_{1}$ vs. $v_{2}$ vs. $v_{3}$ denotes the comparison between the three versions of evaluations.
When human and LLM judges grade answers involving model-generated chain-of-thought rationales, consistently and precisely assigning the same score in 1-10 across multiple judgments is challenging, even if a reference correct answer is provided and each score in 1-10 has a particular and comprehensive description. Meanwhile, it is impossible to cover all situations in the descriptions of scores. Nonetheless, existing LLM-as-a-Judge approaches only adopt prompts with several aspects of constraints, which is undoubtedly very difficult because LLMs do not know exactly what kind of answer corresponds to each score.
Newly expanded context windows of LLMs allow researchers to investigate ICL with more shots than the zero-shot and few-shot regimes, namely many-shot ICL. To fully investigate the many-shot ICL, Agarwal et al. (2024) propose two settings of many-shot ICL (i.e., Reinforced ICL and Unsupervised ICL) to explore scaling in-context learning to hundreds or thousands of in-context examples and find performance improvements across multiple tasks. More importantly, they show that using the many-shot in-context examples with chain-of-thought rationales generated through the zero-shot regime is effective, and the many-shot ICL may overcome the pre-training biases of LLMs, whereas few-shot ICL struggles. Therefore, the intuitive idea is to use the many-shot ICL, allowing LLM-as-a-Judge to see the zero-shot evaluations of similar questions and answers first and then scoring examples before scoring. Therefore, an interesting question arises:
@@
Motivated by prior findings and the above issue, we verify the consistency of the widely used prompts of LLM-as-a-Judge Zheng et al. (2023), as shown in Table 3. Concretely, the consistency experiments are based on the entire test set of GSM8K Cobbe et al. (2021), where the inference answers are obtained based on LLaMA3-70B (The pipeline as illustrated as in Figure 2 and details in §$\S$ 3). Figure 1 presents that the consistency between the two versions of evaluations is low, with nearly half of the ratings being inconsistent.
Inspired by the above phenomena and issues, in this paper, we investigate whether many-shot in-context learning helps long-context LLM judges. Motivated by Agarwal et al. (2024), we introduce two versions of many-shot ICL prompts via Reinforced ICL and Unsupervised ICL for LLM-as-a-Judge. The former utilizes in-context examples with model-generated rationales, and the latter removes rationales used in the former, details in §$\S$ A.1. Meanwhile, we reveal a novel symbol bias in GPT-4o-as-a-Judge and explore a simple approach for mitigating this issue. Experiments show that many-shot ICL can help GPT-4o-as-a-Jduge obtain higher quality and consistency evaluation results. As the number of in-context examples increases, the quality and consistency of evaluation improves significantly. Furthermore, we further verify the effectiveness of the proposed simple yet effective approach for mitigating the symbol bias in pairwise comparison of GPT-4o-as-a-Jduge.
@@
  • Firstly, we propose investigating many-shot ICL to help long-context LLM judges, such as GPT-4o. To the best of our knowledge, we are the first to explore and introduce many-shot in-context learning to assist long-context LLMs as judges in single answer grading of GPT-4o-as-a-Judge and scale the number of in-context examples to verify the effectiveness.
  • Second, inspired by the Counting-Stars benchmark Song et al. (2024), we reveal a novel potential bias in pairwise comparison of GPT-4o-as-a-Judge, namely symbol bias, and propose a simple yet effective approach to mitigate it.
  • Finally, through experimental results, many-shot ICL can help GPT-4o-as-a-Jduge obtain higher quality and consistency evaluation results. Meanwhile, we demonstrate the effectiveness of the symbol bias mitigation method.
@@
Figure 3: Consistency between two versions of judgment results. Concretely, the bar corresponding to "0" on the x-axis represents the number of samples with consistent and inconsistent ratings in comparing evaluation results obtained twice using GPT-4o as the judge in the zero-shot regime. In addition, the zero-shot generated rationales are used for Reinforced ICL and appended to the prompt for Unsupervised ICL. The bar corresponding to "$2^{n}$" on the x-axis represents the consistency of using the GPT-4o as a judge in the many-shot Reinforced ICL regime.
@@ 2.1 Background of Many-Shot ICL
LLMs excel at few-shot in-context learning, which involves learning from a few input-output demonstrations (“shots”) provided in context at inference without weight updates Brown et al. (2020). Newly expanded long-context LLMs allow us to investigate ICL with hundreds of in-context examples Li et al. (2023); Agarwal et al. (2024). Due to some limitations, we only examine GPT-4o-as-a-Judge in a many-shot regime in this paper.
@@ 2.2 Recalling Judgment Biases
LLM judges possess potential biases, which have been widely explored Wang et al. (2023b); Wu and Aji (2023); Zheng et al. (2023); Chen et al. (2024). For example, positional bias refers to the phenomenon where, during pairwise comparisons, LLM judges tend to favor one side of a pair regardless of the actual quality of the answers. In this paper, we investigate leveraging many-shot ICL to help GPT-4o as a better Judge, maybe through reducing judgment biases.
@@ 2.3.1 Single Answer Grading
In this variation, the GPT-4o Judge is required to assign a score with a reason to a single answer directly. Inspired by the recent work Agarwal et al. (2024), we adopt the Reinforced ICL, which uses model-generated rationales as the in-context examples. Meanwhile, we also design the prompt within the Unsupervised ICL. The 4-shot Reinforced ICL and 1-shot Unsupervised ICL append 4-shots with prompts are presented in Table 1 and Table 5.
@@ 2.3.2 Pairwise Comparison
The GPT-4o Judge is presented with a question and two answers and tasked with determining which is better or declaring a tie. The prompt used for this scenario is given in Table 3. In this work, we adopt the pairwise comparison of GPT-4o-as-a-Judge to evaluate the quality of the GPT-4o judgments results in the zero-shot regime.
Figure 4: Consistency between two versions of judgment results.
Figure 5: Consistency between two versions of judgment results.
Figure 6: Consistency between two versions of judgment results.
@@ 3 Experiments
In this section, we introduce the experimental settings, consistency evaluation in single answer grading of GPT-4o-as-a-Jduge, quality evaluation via pairwise comparison of GPT-4o-as-a-Jduge, and reveal the symbol bias in GPT-4o-as-a-Jduge.
@@ 3.1 Experimental Settings
To analyze the GPT-4o judge in the many-shot regime, we use LLaMA3-70B to generate answers for each question in GSM8K Cobbe et al. (2021) with a temperature of 0.7. Details of the used benchmark are provided in §$\S$ A.3 (Appendix). There are two reasons for using LLaMA3 to infer the answers to the problems in GSM8K. First, the quality of answers obtained by LLaMA3 may be high or low, which makes it beneficial to use these answers to analyze the GPT-4o-as-a-Judge. Second, if GPT-4o is adopted to infer the answers, using GPT-4o to judge again has potential judgment biases. We use the training set of GSM8K as the sampling pool of in-context samples and use the first 200 samples of the test set in GSM8K for experiments in this paper. Specifically, we randomly sample in-context examples for each ${K}$-shot prompt in each test data for reliable results. To ensure that using more shots provides additional information, any ${K}$-shot ICL prompt in our experiments includes all in-context examples from prompts with less than ${K}$ examples (all examples in the training set).
Furthermore, inspired by Counting-Stars Song et al. (2024), even if too many shots are provided, the GPT-4o may not be able to utilize all of them. Because, in the Counting-Stars benchmark, when the number of pieces of evidence reaches 32, long-context LLMs (e.g., GPT-4 Turbo, Gemini 1.5 Pro, and Claude3 Opus) may no longer accurately obtain all of them. Therefore, adding more many-shot in-context examples is probably not captured by LLMs for learning as a reference to judge. However, the many-shot regime and Counting-Stars are substantially different, so there is no noise from "haystack". Hence, we set the maximum number of the in-context examples to 128, i.e., 128-shots.
@@
3.2 Consistency Evaluation
We investigate the consistency between different versions of evaluation results generated by the GPT-4o judge. Here, the single answer grading evaluation results can be used to compare the consistency between different versions. As shown in Figure 3, the bar corresponding to "$2^{n}$" on the x-axis represents the number of samples with consistent and inconsistent ratings in comparing evaluation results obtained twice using GPT-4o-as-a-Judge in the many-shot Reinforced ICL regime. From the results, consistency improves as we increase the number of shots provided as in-context examples during inference.
Recent studies Wang et al. (2023b); Zheng et al. (2023); Chen et al. (2024) have shown that the judgment performance of GPT-4 as a judge is highly in agreement with those of humans. However, both human and LLM judgments are subject to potential biases Chen et al. (2024). By analyzing prior studies, we suppose it unnecessary to obtain utterly accurate evaluation results (because this is difficult) by using LLMs as judges. It is only required to ensure that the evaluation results are highly consistent multiple times so that the single answer grading of GPT-4o-as-a-Judge may be effective. From all results in this work, we find that the many-shot ICL examples help the judgment of LLMs more consistently, which is essential. We consider that the main reason may be that the many-shot in-context examples mitigate the potential biases of the GPT-4o judge in the zero-shot regime.
Meanwhile, we also implemented experiments in the many-shot Unsupervised ICL, but the results show that this regime may be unsuitable in the scenario of acting as a judge because the problem of each sample is the same, but the scored questions and answers are different. From the results in Figure 4, Figure 5, and Figure 6, it can be seen that the consistency corresponds to the number of appended in-context samples with model-generated rationales but nearly not to the number of in-context examples without model-generated rationales.
In addition, both the few-shot and many-shot regimes are sensitive to the selection and order of in-context examples. In the experimental setting of this paper, almost no test data will have the same in-context examples. Even under this condition, the evaluation results also show a high consistency, demonstrating the effectiveness of many-shot ICL in helping GPT-4o-as-a-Judge.
@@ 3.3 Quality Evaluation
After the consistency evaluation, an important question arises about: Does a high consistency refer to high-quality judgments? Around this issue, we pair-wise compare model-generated judgment rationales between zero-shot and many-shot regimes using the designed prompts shown in Table 2.
From Figure 7 Compare(A, B), we can find that the evaluations obtained in the many-shot regime are significantly better than those in the zero-shot regime. However, as mentioned before, the positional bias of GPT-4o-as-a-Judge may cause these results, so we performed a second comparison by swapping the positions. As shown in Figure 7 Compare(B, A), we observe that as in-context examples increase, the judgment results in the many-shot regime gradually turn around, that is, a higher winning rate. To obtain fairness results, we integrate the above two results (Compare(A, B) and Compare(B, A)), as shown in Figure 8. It can be seen that after mitigating the positional bias, the judgment quality in the many-shot regime is still better than that in the zero-shot regime.
3.4 Revealing Symbol Bias
Inspired by Song et al. (2024), LLMs revealed a preference for specific symbols under stress testing (details are described in the Appendix). Therefore, an interesting question arises about: Does the GPT-4o prefer to choose the answer with the symbol A rather than B? To verify this conjecture, we swap the answers corresponding to symbols A and B, as shown in Figure 2. From Figure 7 Compare(A, B) and Compare(B, A), it can be seen that the results are different from the above experiments. Actually, the results of Compare(A, B) and Compare(A, B) should be similar, and the results of Compare(B, A) and Compare(B, A) should be similar. This phenomenon shows that symbol bias does exist when adopting GPT-4o-as-a-Judge.
@@ 3.4 Revealing Symbol Bias
Furthermore, there is an interesting hypothesis that the phenomenon of the positional bias we think is not actually caused by the positions of answers, instead of the symbols of answers. In other words, there may be no positional bias in using LLMs as judges in pairwise comparison but rather the symbol bias. The experimental results from Figure 8 Compare(A, B) and Compare(A, B) can demonstrate this hypothesis well.
@@ 4 Related Work
LLMs have exhibited remarkable general generation capabilities, positioning themselves as powerful assistants Zhao et al. (2023); OpenAI (2023). With the rapid progression of LLMs, evaluating their proficiency in adhering to human instructions is imperative. Given the advanced capabilities of LLMs, researchers have begun adopting these models to judge the performance of LLMs in following human instructions Koo et al. (2023); Liusie et al. (2023); Liu et al. (2023); Zhu et al. (2023); Lu et al. (2023); Fu et al. (2023b); Zheng et al. (2023); Wang et al. (2023b); Chen et al. (2024). Notably, the evaluation paradigm introduced by Zheng et al. (2023) has gained widespread adoption. However, LLMs as judges are revealed to have potential biases Wang et al. (2023b); Chen et al. (2024), which leads to higher uncertainty and inconsistency during the evaluation using LLMs, questioning the validity of LLM-as-a-Judge.